# Gaitsgory’s central functor and the Arkhipov–Bezrukavnikov equivalence in mixed characteristic

### v3

@@ 1.4. Acknowledgements
This paper owes a lot to the machinery developed in [AGLR22], so we are first of all thankful to Ian Gleason and Timo Richarz for their contribution to that project, and also further conversations involving this project. We benefited from a workshop on the Bezrukavnikov equivalence in Essen–Münster–Bonn organized by the first two authors together with Ulrich Görtz, Eugen Hellmann, and Konrad Zou, and we are thankful to all of the participants. This project has received funding from the Excellence Cluster of the Universität Münster, and the ERC Consolidator Grant 770936 via Eva Viehmann. The fourth author benefited from participating in the special academic year at the Institute for Advanced Study. He thanks Zurich Insurance Company for the support and organizers and all participants of the special year program. Part of the project was conducted when the fourth author was visiting the Max-Planck Institute for Mathematics, he is grateful to the institute for host. We would also like to acknowledge the relevant discussions we had or feedback we got from Pramod Achar, Robert Cass, Harrison Chen, Jean-François Dat, Arnaud Étève, Thomas Haines, Linus Hamann, David Hansen, Xuhua He, Eugen Hellmann, Claudius Heyer, Tasho Kaletha, Arthur-César Le Bras, Lucas Mann, Cédric Pépin, Simon Riche, Peter Scholze, Thibaud van den Hove, Eva Viehmann, Yifei Zhao, Xinwen Zhu, Konrad Zou.
@@ 1.5. Notation
Unfortunately, we will have to use a lot of notations. Thus, let’s get over this and define the following objects, which will occur in the whole text.
@@
  • $F$ a $p$-adic field with ring of integers $O$, and (perfect) residue field $k$.
@@
  • $\bar{{\mathbb{X}}}_{\bullet}:={\mathbb{X}}_{\bullet,\breve{\Gamma}}$, where the subscript $\breve{\Gamma}$ denotes the coinvariants.
@@
  • ${\mathcal{G}}/O$ the parahoric model of $G$ associated with the facet ${\mathbf{f}}$
@@ 2.1. Schubert varieties and convolution
of morphisms of ind-perfect schemes. The first morphism is affine as $X\subset{\mathrm{Fl}}_{{\mathcal{G}}}$ is a closed immersion, and the second is affine by Lemma 2.5. If ${\mathcal{G}}$ is Iwahori, then the affineness of $m\colon X\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w}\to{\mathrm{Fl}}_{{\mathcal{I}}}$ can be checked after passing to the quotient Hecke stacks
$[L^{+}{\mathcal{I}}\backslash X\times^{L^{+}{\mathcal{I}}}{\mathrm{Fl}}_{{\mathcal{I}},w}]\to{\operatorname{Hk}}_{{\mathcal{I}}}$ (2.19)
Forgetting about $L^{+}{\mathcal{I}}$-equivariance on the right, we reduce to the previous case. ∎
@@
$m(X\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w})\subset\cup_{x\in S_{X}}{\mathrm{Fl}}_{{\mathcal{I}},xw},\quad m({\mathrm{Fl}}_{{\mathcal{I}},w}\tilde{\times}X)\subset\cup_{x\in S_{X}}{\mathrm{Fl}}_{{\mathcal{I}},wx}.$
@@
$m({\mathrm{Fl}}_{{\mathcal{I}},w_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}})\subset\cup_{x\in S_{X_{1}}}{\mathrm{Fl}}_{{\mathcal{I}},xw^{\prime}},\quad m({\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w_{2}})\subset\cup_{x\in S_{X_{2}}}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}x},$
@@
${\mathrm{Fl}}_{{\mathcal{I}},s_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\subset{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\sqcup{\mathrm{Fl}}_{{\mathcal{I}},s_{1}w^{\prime\prime}},\quad{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},s_{2}}\subset{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\sqcup{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}s_{2}},$
@@
${\mathrm{Fl}}_{{\mathcal{I}},w_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},s_{1}}\simeq{\mathrm{Fl}}_{{\mathcal{I}},s_{2}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w_{2}}\simeq{\mathrm{Fl}}_{{\mathcal{I}},w}$
@@ 2.2. Constant terms and semi-infinite orbits
Next we describe the closure relations for the stratification of ${\mathrm{Fl}}_{{\mathcal{I}}}$ by the ${\mathcal{S}}_{w},w\in W$. As this is a geometric questions, we may assume $F=\breve{F}$ for this. Then we have to define the dominant cocharacters $\bar{{\mathbb{X}}}_{\bullet}^{+}$ in $\bar{{\mathbb{X}}}_{\bullet}$. Recall that ${\mathbb{X}}_{\bullet}(S)^{+}$ denotes the ($B$-)dominant cocharacters for $S$. As we assumed $F=\breve{F}$, we get that
@@ 3. Cohomology of the affine flag variety
be the full subcategory consisting of objects with perfect stalks. Let $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ be the Grothendieck group of ${\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$. Since points in ${\mathrm{Fl}}_{{\mathcal{G}}}$ have connected stabilizers under the $L^{+}{\mathcal{I}}$-action, $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ identifies with the Grothendieck group of the category of $L^{+}{\mathcal{I}}$-equivariant sheaves on ${\mathrm{Fl}}_{{\mathcal{G}}}$. Consequently, $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ is a free abelian group on the classes of the intersection complexes ${\operatorname{IC}}_{({\mathcal{I}},{\mathcal{G}}),w}=j_{!\ast}(\Lambda)$ of ${\mathrm{Fl}}_{({\mathcal{I}},{\mathcal{G}}),\leq w}$.
@@
Note that by definition ${\mathscr{D}}_{\mathrm{cons}}^{\leq 0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ (resp. ${\mathscr{D}}_{\mathrm{cons}}^{\geq 0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$) is spanned by the non-negative (resp. non-positive) shifts of the $\Delta_{({\mathcal{I}},{\mathcal{G}}),v}$ (resp. $\nabla_{({\mathcal{I}},{\mathcal{G}}),v}$) for $v\in W/W_{\mathbf{f}}$. By Lemma 2.6, the convolution map $m\colon{\mathrm{Fl}}_{{\mathcal{I}},w}\tilde{\times}{\mathrm{Fl}}_{({\mathcal{I}},{\mathcal{G}}),\leq v}\rightarrow{\mathrm{Fl}}_{\mathcal{I}}$ is affine. Now,
$\Delta_{{\mathcal{I}},w}\ast\nabla_{({\mathcal{I}},{\mathcal{G}}),v}=Rm_{!}({\Lambda}\tilde{\boxtimes}\nabla_{v})$ (3.11)
@@
$\nabla_{{\mathcal{I}},w}\ast\Delta_{({\mathcal{I}},{\mathcal{G}}),v}=Rm_{\ast}(\Delta_{w}\tilde{\boxtimes}{\Lambda})$ (3.12)
@@
    $\operatorname{supp}(\Delta_{w}*{\mathcal{F}})\subseteq w\cdot A_{{\mathcal{F}}},\quad\operatorname{cosupp}(\nabla_{w}*{\mathcal{F}})\subseteq w\cdot A_{{\mathcal{F}}}$,
@@
    $\operatorname{supp}({\mathcal{F}}*\Delta_{w})\subseteq A_{{\mathcal{F}}}\cdot w,\quad\operatorname{cosupp}({\mathcal{F}}\ast\nabla_{w})\subseteq A_{\mathcal{F}}\cdot w$.
@@ 3.2. Wakimoto sheaves
Note that all the transition morphisms in the colimit are isomorphisms. In particular, we can conclude (by invertibility of $\nabla_{t_{\bar{\nu_{2}}}})$, cf. Lemma  3.3) that
@@
The first statement follows from 3.5, the second from 3.6 and the third is implicit in the discussion of monoidality of ${\mathscr{I}}$. Let us discuss the forth statement. Using (3), the invertibity of ${\mathscr{I}}_{\bar{\nu}}(\Lambda)$ and the definition of the semi-infinite Bruhat order $\preceq$ reduces by suitable convolution to the case that $\bar{\nu},\bar{\mu}\in\bar{{\mathbb{X}}}_{\bullet}$ are dominant. Then ${\mathscr{I}}_{\bar{\mu}}(M)=\nabla_{t_{\bar{\mu}}}(M),{\mathscr{I}}_{\bar{\nu}}(N)=\nabla_{t_{\bar{\nu}}}(N)$ and thus by 2.2
@@
Define the Wakimoto category as the full subcategory $\operatorname{Wak}:=\langle{\mathscr{I}}\rangle$ of ${\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{\mathcal{I}})$ generated by the essential image of ${\mathscr{I}}$ under cones. An object ${\mathcal{F}}\in\operatorname{Ob}(\operatorname{Wak})$ is called Wakimoto filtered.
@@
Given ${\mathcal{F}}\in\operatorname{Ob}(\operatorname{Wak})$, we show the existence of a final morphism ${\mathcal{F}}_{\Omega}\to{\mathcal{F}}$, such that ${\mathcal{F}}_{\Omega}\in\operatorname{Ob}(\operatorname{Wak}_{\Omega})$. By induction on the length of the filtration, we can write ${\mathcal{F}}$ as an extension of ${\mathcal{G}}={\mathcal{G}}_{\Omega}$ by ${\mathscr{I}}_{\bar{\nu}}(M)$ for some $\bar{\nu}\in\bar{\mathbb{X}}_{\bullet}$ and $M\in{\mathscr{D}}_{\mathrm{\acute{e}t}}^{b}({\operatorname{Spec}}\,k)$. If $\bar{\nu}\in\Omega$, we are done. Otherwise, we have $\bar{\mu}\not\preceq\bar{\mu}$ for all $\mu\in\Omega$ by the lower set hypothesis on $\Omega$, so we get $R{\operatorname{Hom}}({\mathscr{I}}_{\bar{\nu}}(M),{\mathcal{G}})=0$ by Lemma 3.12 and hence there exists a (unique) splitting ${\mathcal{F}}\simeq{\mathscr{I}}_{\bar{\nu}}(M)\oplus{\mathcal{G}}$. One concludes that ${\mathcal{F}}_{\Omega}$ exists and identifies with ${\mathcal{G}}$ via the given map to ${\mathcal{F}}$ (again by Lemma 3.12). ∎

### v4

@@ 1.4. Acknowledgements
This paper owes a lot to the machinery developed in [AGLR22], so we are first of all thankful to Ian Gleason and Timo Richarz for their contribution to that project, and also further conversations involving this project. We benefited from a workshop on the Bezrukavnikov equivalence in Essen–Münster–Bonn organized by the first two authors together with Ulrich Görtz, Eugen Hellmann, and Konrad Zou, and we are thankful to all of the participants. This project has received funding from the Excellence Cluster of the Universität Münster, and the ERC Consolidator Grant 770936 via Eva Viehmann. The fourth author benefited from participating in the special academic year at the Institute for Advanced Study. He thanks Zurich Insurance Company for the support and organizers and all participants of the special year program. Part of the project was conducted when the fourth author was visiting the Max-Planck Institute for Mathematics, he is grateful to the institute for host. We would also like to acknowledge the relevant discussions we had or feedback we got from Pramod Achar, Robert Cass, Harrison Chen, Jean-François Dat, Gurbir Dhillon, Arnaud Étève, Thomas Haines, Linus Hamann, David Hansen, Xuhua He, Eugen Hellmann, Claudius Heyer, Tasho Kaletha, Arthur-César Le Bras, Lucas Mann, Cédric Pépin, Simon Riche, Peter Scholze, Jeremy Taylor, Thibaud van den Hove, Eva Viehmann, Yifei Zhao, Xinwen Zhu, Konrad Zou. We also thank the anonymous referee for carefully reading the paper and suggesting several improvements.
@@ 1.5. Notation
Unfortunately, we will have to use a lot of notations. Thus, let us address this once and for all and define the following objects, which will occur in the whole text.
@@
  • $F$ a finite extension of either $\mathbb{Q}_{p}$ or ˘$\breve{\mathbb{Q}}_{p}$ with ring of integers $O$, and (perfect) residue field $k$.
@@
  • $\bar{{\mathbb{X}}}_{\bullet}:={\mathbb{X}}_{\bullet,\breve{\Gamma}}$, where the subscript $\breve{\Gamma}$ denotes the coinvariants,
@@
  • ${\mathcal{G}}/O$ the parahoric model of $G$ associated with the facet ${\mathbf{f}}$,
@@ 2.1. Schubert varieties and convolution
of morphisms of ind-perfect schemes. The first morphism is affine as $X\subset{\mathrm{Fl}}_{{\mathcal{G}}}$ is a closed immersion, and the second is affine by Lemma 2.5.
If ${\mathcal{G}}$ is Iwahori, then the affineness of $m\colon X\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w}\to{\mathrm{Fl}}_{{\mathcal{I}}}$ can be checked after passing to the quotient Hecke stacks
$[L^{+}{\mathcal{I}}\backslash\tilde{X}\times^{L^{+}{\mathcal{I}}}{\mathrm{Fl}}_{{\mathcal{I}},w}]\to{\operatorname{Hk}}_{{\mathcal{I}}}$ (2.19)
where $\tilde{X}\subset LG$ is the natural $L^{+}{\mathcal{I}}$-torsor over $X\subset{\mathrm{Fl}}_{{\mathcal{I}}}$. Forgetting about $L^{+}{\mathcal{I}}$-equivariance on the right and applying the inverse map of $LG$ and its subgroup $L^{+}{\mathcal{I}}$, we arrive at the convolution map
$Y\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w^{-1}}\to{\mathrm{Fl}}_{{\mathcal{I}}}$ (2.20)
where $Y=\tilde{Y}/L^{+}{\mathcal{I}}$ and $\tilde{Y}=\tilde{X}^{-1}\subset LG$. It suffices to prove this map is affine, but we already treated it in the first paragraph. ∎
@@
$m(X\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w})\subset\cup_{x\in S_{X}}{\mathrm{Fl}}_{{\mathcal{I}},xw},\quad m({\mathrm{Fl}}_{{\mathcal{I}},w}\tilde{\times}X)\subset\cup_{x\in S_{X}}{\mathrm{Fl}}_{{\mathcal{I}},wx}.$ (2.21)
@@
$m({\mathrm{Fl}}_{{\mathcal{I}},w_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}})\subset\cup_{x\in S_{X_{1}}}{\mathrm{Fl}}_{{\mathcal{I}},xw^{\prime}},\quad m({\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w_{2}})\subset\cup_{x\in S_{X_{2}}}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime}x},$ (2.22)
@@
$m({\mathrm{Fl}}_{{\mathcal{I}},s_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}})\subset{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\sqcup{\mathrm{Fl}}_{{\mathcal{I}},s_{1}w^{\prime\prime}},\quad m({\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},s_{2}})\subset{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}}\sqcup{\mathrm{Fl}}_{{\mathcal{I}},w^{\prime\prime}s_{2}},$ (2.23)
@@
${\mathrm{Fl}}_{{\mathcal{I}},w_{1}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},s_{1}}\simeq{\mathrm{Fl}}_{{\mathcal{I}},s_{2}}\tilde{\times}{\mathrm{Fl}}_{{\mathcal{I}},w_{2}}\simeq{\mathrm{Fl}}_{{\mathcal{I}},w}$ (2.24)
@@ 2.2. Constant terms and semi-infinite orbits
Next we describe the closure relations for the stratification of ${\mathrm{Fl}}_{{\mathcal{I}}}$ by the ${\mathcal{S}}_{w},w\in W$. As this is a geometric question, we may assume $F=\breve{F}$ for this. Then we have to define the dominant cocharacters $\bar{{\mathbb{X}}}_{\bullet}^{+}$ in $\bar{{\mathbb{X}}}_{\bullet}$. Recall that ${\mathbb{X}}_{\bullet}(S)^{+}$ denotes the ($B$-)dominant cocharacters for $S$. As we assumed $F=\breve{F}$, we get that
@@ 3. Cohomology of the affine flag variety
be the full subcategory consisting of objects with perfect stalks. Let $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ be the Grothendieck group of ${\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$. Since points in ${\mathrm{Fl}}_{{\mathcal{G}}}$ have connected stabilizers under the $L^{+}{\mathcal{I}}$-action, $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ identifies with the Grothendieck group of the category of $L^{+}{\mathcal{I}}$-equivariant sheaves on ${\mathrm{Fl}}_{{\mathcal{G}}}$. Consequently, $K_{0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ is a free abelian group on the classes of the intersection complexes ${\operatorname{IC}}_{({\mathcal{I}},{\mathcal{G}}),w}:=j_{w,!\ast}(\Lambda)$ of ${\mathrm{Fl}}_{({\mathcal{I}},{\mathcal{G}}),\leq w}$.
@@
We follow [AR, Lemma 4.1.7]. Note that by definition ${\mathscr{D}}_{\mathrm{cons}}^{\leq 0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$ (resp. ${\mathscr{D}}_{\mathrm{cons}}^{\geq 0}({\operatorname{Hk}}_{({\mathcal{I}},{\mathcal{G}})})$) is spanned by the non-negative (resp. non-positive) shifts of the $\Delta_{({\mathcal{I}},{\mathcal{G}}),v}$ (resp. $\nabla_{({\mathcal{I}},{\mathcal{G}}),v}$) for $v\in W/W_{\mathbf{f}}$. By Lemma 2.6, the convolution map $m\colon{\mathrm{Fl}}_{{\mathcal{I}},w}\tilde{\times}{\mathrm{Fl}}_{({\mathcal{I}},{\mathcal{G}}),\leq v}\rightarrow{\mathrm{Fl}}_{\mathcal{I}}$ is affine. Now,
$\Delta_{{\mathcal{I}},w}\ast\nabla_{({\mathcal{I}},{\mathcal{G}}),v}=Rm_{!}({\Lambda}\tilde{\boxtimes}\nabla_{({\mathcal{I}},{\mathcal{G}}),v})$ (3.11)
@@
$\nabla_{{\mathcal{I}},w}\ast\Delta_{({\mathcal{I}},{\mathcal{G}}),v}=Rm_{\ast}(\Delta_{{\mathcal{I}},w}\tilde{\boxtimes}{\Lambda})$ (3.12)
@@
    $\operatorname{supp}(\Delta_{w}*{\mathcal{F}})\subseteq w\cdot A_{{\mathcal{F}}}$, and dually $\operatorname{cosupp}(\nabla_{w}*{\mathcal{F}})\subseteq w\cdot A_{{\mathcal{F}}}$,
@@
    $\operatorname{supp}({\mathcal{F}}*\Delta_{w})\subseteq A_{{\mathcal{F}}}\cdot w$, and dually $\operatorname{cosupp}({\mathcal{F}}\ast\nabla_{w})\subseteq A_{\mathcal{F}}\cdot w$.
@@ 3.2. Wakimoto sheaves
Note that all the transition morphisms in the colimit are isomorphisms. In particular, we can conclude (by invertibility of $\nabla_{t_{\bar{\nu_{2}}}}$, cf. Lemma 3.3) that
@@
The first statement follows from 3.5, the second from 3.6 and the third is implicit in the discussion of monoidality of ${\mathscr{I}}$. Let us discuss the fourth statement. Using (3), the invertibity of ${\mathscr{I}}_{\bar{\nu}}(\Lambda)$ and the definition of the semi-infinite Bruhat order $\preceq$ reduces by suitable convolution to the case that $\bar{\nu},\bar{\mu}\in\bar{{\mathbb{X}}}_{\bullet}$ are dominant. Then ${\mathscr{I}}_{\bar{\mu}}(M)=\nabla_{t_{\bar{\mu}}}(M),{\mathscr{I}}_{\bar{\nu}}(N)=\nabla_{t_{\bar{\nu}}}(N)$ and thus by 2.2
@@
Define the Wakimoto category as the full subcategory $\operatorname{Wak}:=\langle{\mathscr{I}}\rangle$ of ${\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{\mathcal{I}})$ generated by the essential image of ${\mathscr{I}}$ under cones and shifts. An object ${\mathcal{F}}\in\operatorname{Ob}(\operatorname{Wak})$ is called Wakimoto filtered.
@@
Given ${\mathcal{F}}\in\operatorname{Ob}(\operatorname{Wak})$, we show the existence of a final morphism ${\mathcal{F}}_{\Omega}\to{\mathcal{F}}$, such that ${\mathcal{F}}_{\Omega}\in\operatorname{Ob}(\operatorname{Wak}_{\Omega})$. By induction on the length of the filtration, we can write ${\mathcal{F}}$ as an extension of ${\mathcal{G}}={\mathcal{G}}_{\Omega}$ by ${\mathscr{I}}_{\bar{\nu}}(M)$ for some $\bar{\nu}\in\bar{\mathbb{X}}_{\bullet}$ and $M\in{\mathscr{D}}_{\mathrm{cons}}([*/S_{k}])$. If $\bar{\nu}\in\Omega$, we are done. Otherwise, we have $\bar{\mu}\not\preceq\bar{\mu}$ for all $\mu\in\Omega$ by the lower set hypothesis on $\Omega$, so we get $R{\operatorname{Hom}}({\mathscr{I}}_{\bar{\nu}}(M),{\mathcal{G}})=0$ by Lemma 3.12 and hence there exists a (unique) splitting ${\mathcal{F}}\simeq{\mathscr{I}}_{\bar{\nu}}(M)\oplus{\mathcal{G}}$. One concludes that ${\mathcal{F}}_{\Omega}$ exists and identifies with ${\mathcal{G}}$ via the given map to ${\mathcal{F}}$ (again by Lemma 3.12). ∎